OpenStudy (anonymous):

What is the fifth term of the sequence whose first term is 10 and whose common ratio is 1/2 ?

OpenStudy (anonymous):

\[10\times (\frac{1}{2})^4\]

OpenStudy (anonymous):

or since \(2^4=16\) it is \(\frac{10}{16}\) maybe can reduce

OpenStudy (anonymous):

I got 50

OpenStudy (anonymous):

another way first term is 10 second term is 5 third term is 2.5 fourth term is 1.25 fifth term is .625

OpenStudy (anonymous):

oh no, it is not \(10\times 2^4\) it is \(\frac{10}{2^4}\)

OpenStudy (anonymous):

thanks